OpenStudy (anonymous):

Please help with two questions? Simplify each product. Write in standard form.

OpenStudy (anonymous):

\[q(11+8q-2q ^{2})\]

OpenStudy (anonymous):

\[(4a-7)(8a+3)\]

OpenStudy (anonymous):

Hello?

OpenStudy (ghostgate):

First of all nice profile picture! Secondly lets try to deduce this problem. As far as I know from Being up in Geometry currently I would suggest you would do: q(11 + 8q - 2q^2) Well this is q x everything in there thanks to the rules of Algebra. So it will look like: 11q + 8q^2 - 2q^3 Now you got to reduce it even further: Add like terms so: 11q + 8q^2 = 19q^3 and: 19q^3 - 2q^3 = 17q - That should be your reduced answer if not than I'll have to recheck and see what your trying to find. Now since you have no equal sign I can't reduce it any futher otherwise it would be the (= # divided by 17 and than q = #) Now for the second one I forgot. I've done them before except my brain just can't think right now. Sorry man. I may be incorrect on the first one, but I'm not sure. Anyways I hope that helps! G'day!

OpenStudy (anonymous):

So is 19q^3 - 2q^3 = 17q the answer you got then?

OpenStudy (ghostgate):

If my math is correct yes it is. If not than I don't know how exactly I got it than. I believe the lowest reduced answer is 17q.

OpenStudy (anonymous):

Ok. Thanks for explaining a bit then.

OpenStudy (ghostgate):

No problem.